# Tailoring transport in quantum spin chains via disorder and collisions

### v3

@@ Abstract
We systematically investigate the interplay of disorder and space-time heterogeneous collisional noise in shaping the transport dynamics of an anisotropic XXZ spin chain. Using stochastic collision models to simulate interaction with the environment, we explore the localization-delocalization transitions across regimes with single and multiple excitations. We find that space homogeneous collisions occurring at low rates favor the shaping of regions where the localization degree sets in the form of subsequent plateaus at a rate and duration universally scaling with the collision rate. We also find that interactions among the excitations favor this process even for tiniest levels of disorder.
@@ I Introduction
Quantum transport governs charge, spin and energy transfer in a wide variety of systems, from solid-state physics and spintronics to quantum technologies [1, 2, 3, 4]. In fact, by exploring the effects of disorder, noise, and interactions, we can gain insights into the mechanisms driving localization-delocalization transitions, thermalization, and the emergence of collective phenomena in low-dimensional and complex systems. In particular, current technologies enable the study of quantum transport with the engineering of reduced-dimensionality systems in solid state and in AMO (atomic, molecular, and optical) platforms [5, 6].
@@
Paradigmatic many-body Hamiltonians in open quantum networks are important models as they represent essential functions and can be engineered in tunable current experiments used as quantum simulators [7]. We are interested in biology applications [8, 9] and on coarse-grained models within the so-called quantum-like paradigm [10, 11]. In general, regardless of the presence of quantum effects in the performance of specific biological processes, the statistics of quantum mechaniscs and the open quantum system framework [12, 13, 14, 15] provides us with a well-suited mathematical toolbox to describe complex phenomena, both quantum and classical, into potentially simpler and representative models [16, 11].
@@
Quantum effects in biological systems present a fascinating yet debated area of research, challenging conventional views on the role of quantum mechanics in the warm and noisy environment of living organisms. These physics-based models can incorporate quantum effects into the description of microscopic phenomena allowing us to study how a noisy and out-of-equilibrium biological environment is influenced by coherent and dissipative quantum effects [8, 17, 18].
@@
In general, the behavior of transport in a quantum system serves as a powerful diagnostic tool for understanding thermalization in disordered conditions [19]. In the absence of interactions, disorder can induce Anderson localization [20], where transport is suppressed due to interference effects, preventing the system from reaching thermal equilibrium. Interactions or noise disrupts this localization mechanism, leading to a transition toward diffusive or even ballistic transport, depending on the system’s parameters [21, 22]. The transition from localized to delocalized regimes provides insight into how thermalization emerges in complex systems. Consequently, disorder is a mechanism that suppresses transport in isolated systems and, in open systems, also a factor that competes with noise to dictate the system’s relaxation dynamics.
@@
In [16], we explored how dissipation modulates transport in an open quantum system using stochastic collision models accounting for space and time heterogeneous collisions. In so doing, we could focus on the interplay between noise, interactions, and coherence under sufficiently general conditions. By tuning noise parameters, we demonstrated controllable modifications to transport properties, identifying regimes where increasing the collision rate enhances transport. Thus, we showed that engineered dissipation can counteract localization effects.
@@
We now focus on the combination of collisional noise and disorder which provides a model that closely mirrors the dynamics of transport in many complex media including those relevant to biological systems. Collisional models replicate approximate interactions with noisy warm environment, while disorder introduces local spatial inhomogeneities. Together, these elements capture key features of energy transfer in systems where coherent and dissipative processes coexist relevant in several fields of physics, like charging and discharging performances in open quantum system batteries [23, 24], transport in an engineered environment of complex systems [25] along with many body localization in 1-D systems [26]. Moreover, the tunability of collisional noise parameters enables the simulation of biologically relevant scenarios, such as the exploitation of noise-assisted transport [27, 28].
@@
In this paper, we investigate the transport properties and thermalization dynamics of a noisy, disordered XXZ spin chain [29, 30, 31, 32], modeled through stochastic collision processes. Among different possibilities, we adopt this many-body model for its capability of describing in a general manner tunneling-based transport of excitations in the presence of coherent anisotropic interactions. We identify localization/delocalization transitions and thus transport windows for multiple excitations tuning the interplay of disorder, anisotropy, and noise.
@@
The paper is organized as follows. In Section II we present the model of our disordered spin-chain interacting with the environment via stochastic collision models and introduce the figures of merit adopted to study transport properties. In Section III, we present the results of our study comparing the cases of one, two and multiple injected excitations also highlighting the entanglement properties in those regimes. Finally, in Section IV we present our conclusions and discuss future perpectives in the light of technological progress in quantum science and biology.
@@ II Model and methods
The conceptual model setting is sketched in Fig. 1 (a). We investigate an integrable generalization of the Heisenberg spin chain that accounts for uni-axial anisotropy [33] with the addition of a disorder term:
@@
$+h\sum_{i=1}^{N}r_{v}(1,i)\sigma^{z}_{i},$ (1)
and subjected to dissipation through noise in the form of stochastic collision models, as depicted in Sec. II.1, with open boundary conditions (OBC). In Eq. 1, $\sigma^{x/y/z}_{i}$ are the Pauli operators acting on spin located at site $i$; $J$ is the spin exchange rate, that can also be seen as the tunneling amplitude for excitation quasiparticles throughout the system; $\Delta$ governs the anisotropy in the form of a spin-spin interaction along the z-axis. In our analysis, we consider the regime with $J=1(>0)$ and then scale every other energy to $J$. As a result, the ground state with disorder $h=0$ is a ferromagnet for $\Delta>1$, becoming a critical paramagnet for $\Delta\leq 1$.
Figure 1: Sketch of our model. (a) Open quantum spin chain with energy levels affected by disorder in the $[-h,h]$ range. Excitations are depicted as flipped spins (green), their dynamics being affected by the tunneling rate $J$ (black arrows), interaction strength $\Delta$ (blue ovals) and the addition of a noisy environment. This is represented by the auxiliary light-brown qubits colliding with the sites over time. (b)-(c) Histograms displaying the number of collisions oundergone by the auxiliary qubits (log scale) against our spins over time tJ for a noise shape parameter of $\nu=100$ in the time-uniform regime, and collision rates $r_{c}=1$ (b) and (c) $r_{c}=0.1$. We use the same simulation parameter: number of sites in the spin is $N=41$ sites, $M=500$ trajectories and final time of simulation $tJ=30$.
The introduction of disorder in the form of a random local field proportional to $\sigma_{i}^{z}$ with amplitude $h$ breaks the integrability of the model. Thus, different ground states would occour under different disorder realizations. The disorder for every spin is then randomly generated, according to the random vector $r_{v}(max,i)$ varying between (0,1) for each site independently, in each individual simulation for every site in a range $[-h,h]$.
When considering the time evolution of excitations or quench dynamics, the disorder term will control different transport regimes, ranging from ballistic to diffusive, and it can lead to the emergence of Many-Body localization (MBL) [34]. In fact, disorder can lead to the breakdown of thermalization, meaning that the system fails to reach thermal equilibrium at long times resulting in transport suppression. Tuning the disorder through $h$ and analyzing the resulting transport behavior, it is possible to map the parameter regimes where localization or delocalization dominates.
@@ II.1 Noise modeling
So far we have discussed how the coherent and disorder-driven terms of our model are described. Here, we introduce dissipation through a particular discretization of the environment degrees of freedom (d.o.f.) that lies in the family of stochastic collision models [9, 35]. This type of environment description comprises heterogeneous interactions in time and space with a bath composed of a set of itinerant d.o.f. This modeling represents an effective approximation to what happens in a biological system [9, 36]. As depicted in Fig.1, the environment d.o.f are modeled as auxiliary qubits (light brown circles) that collide over time with the spins in the chain.
In this approach, we can govern the noise regimes via appropriate distributions describing the probability of system-environment interaction events. This results into a flexible environment description, including non-trivial temporal correlations even remaining within a Markovian-like description. As it is the case for many dissipative dynamical maps, stochastic collisional models are compatible with many local quantum channels. In our case, we account for density-density collisions decorrelating spins from the coherent evolution, thus effectively producing dephasing in the system degree of freedom.
We choose to describe the collisional rate between spins and auxiliary d.o.f. by adopting the especially flexible Weibull distribution [37, 9]:
@@
[…]
When collisions occur according to the probability distribution, a quantum channel is applied [9, 38]. This is given by:
@@
with $U_{\textrm{coll}}=\exp{-i(\pi/2)\sigma_{a}^{x}\otimes\sigma^{z}_{i}}$ representing the collision event and $t_{i}$ being the time before the given collision. For a detailed description, we refer to our previous work [16].
@@ II.2 Characterizing transport
Having described our model, we now introduce the figures of merit that we use to characterize the system transport and quantify the magnetization spreading (delocalization behavior of the excitations). For one excitation, we use the inverse participation ratio (IPR), $\mathrm{IPR}=\sum_{i=1}^{N}\bra{i}\rho(t)\ket{i}^{2},$ where $i$ represents the site index and |$\ket{i}$ represents the single-excitation localized states {|$\Big{\{}\ket{i}\>:\>\ket{i}=\sigma_{i}^{\dagger}\ket{0}^{\otimes N}\Big{\}}$ [9]. The IPR is a measure of localization, bounded between the complete delocalization asymptotic value $\mathrm{IPR}=1/N$, and $\mathrm{IPR}=1$, corresponding to complete localization, i.e. when the excitation remains on a particular site of the network [39].The larger the $\mathrm{IPR}$, the more localized the excitation is over the lattice.
@@
Therefore, to characterize the emergence of parameter regions where the system (de)localizes, we also consider the behaviour of the entanglement entropy for selected scenarios. The entanglement entropy constitutes one of the well-known heralds of MBL [21, 41], as it presents logarithmic growth in many body localized systems. While not generally accessible in experiments with a few important exceptions [42, 43], the study of entanglement entropy in thermal and non-thermal many-body states [44] has helped understand and quantify the degree of correlation and information spreading between system parts, whereas its behaviour can become non-trivial in regions close to the transition point [45]. The entanglement entropy can be characterized via the Von Neumann entropy [46] defined as follows: consider a bi-partite system composed by $A$ and $B$ subsystems, then the entanglement entropy is $S_{vN}=\rho_{A}\log\rho_{A}$, with $\rho_{A}=Tr_{B}\rho$ the reduced density matrix of subsystem $A$. Since we are interested in regions in which transport can be modulated to occur only in well-defined time windows, we analyze the behaviour of entanglement entropy as a complementary figure of merit.
@@ III.1 One excitation case
We begin by building our intuition on the interplay between disorder and noise with the case of one single excitation propagating through the disordered spin chain. In this analysis we highlight the fundamental differences between Anderson localization [21] and noise-induced transport. Anderson localization, driven exclusively by disorder, is characterized by the suppression of transport due to interference effects in the absence of any environmental interactions. In contrast, transport induced by collisional noise introduces a competing mechanism, where dephasing disrupts localization and promotes delocalization over time. By focusing on the one excitation case, we simplify the system dynamics, isolating the interplay between disorder and noise while avoiding complications arising from multiple excitation interactions.
We summarize in Fig. 2 the IPR behavior. Fig. 2(a) summarizes the long-time behavior of the Inverse Participation Ratio (IPR) at time $tJ=30$ as a function of the shape parameter $\nu$ and the collision rate $r_{c}$. We see that for low values of $r_{c}$, the system exhibits higher localization (larger IPR). This occurs because the low collision rate allows for transient localization before delocalization takes place due to noise. More frequent collisions (larger $r_{c}$) disrupt these localized states and drive the system toward delocalization, with the IPR approaching its asymptotic value $1/N$. To see the competing effect of disorder, we show in Fig. 2(b)-(c) the IPR evolution at fixed shape parameter $\nu=100$ in the regime of time-homogeneous collisions. The IPR time evolution is displayed in Fig. 2(b) for various collision rates $r_{c}$ and fixed large disorder range $h=10$. We then show in Fig. 2(c) the IPR versus time for different disorder range constants $h$ from 0.1 to 10 at fixed low collision rate $r_{c}=0.1$. We observe that low collision rates ($r_{c}<1$) and high levels of disorder ($h\gg 1$) lead to transient localization, as evidenced by the formation of sharp plateaus in the IPR. These interesting plateau features reflect moments of temporary stabilization in the system’s dynamics, after which collisions induce delocalization. Conversely, at higher collision rates ($r_{c}\geq 1$), the increased frequency of noise events prevents significant localization, driving the system toward a diffusive regime. Here, the IPR evolves smoothly over time without plateaus formation. The case of $\Delta\neq 0$ was not reported here because the inclusion of an anisotropic interaction term is seen to not qualitatively alter the system’s behavior in the single-excitation regime. For more details see [16].
Figure 2: (De)localization behaviour: one excitation case. Inverse Participation Ratio (IPR) for a spin chain with $N=41$ sites. (a) IPR at long time ($tJ=30$), for $h=10$ and different shape parameters $\nu$ and collision rates $r_{c}$. (b) IPR vs time for large disorder range $h=10$ for different collision rates $r_{c}$. Curves with progressively increasing tickness and color scale refer to increasing values of $r_{c}=0,0.1,0.5,1,5,50,100$ (c) IPR vs time $tJ$ for low $r_{c}=0.1$ and different disorder range $h$. Notice the emergence of plateaus of temporary stabilization of localized behavior. For (b) and (c) the shape parameter is $\nu=100$ in the time-homogeneous collision regime. The data refer to the case of no anisotropy ($\Delta=0$ in Eq. 1) since they are seen to be not significantly altered by finite $\Delta\neq 0$. The simulations were performed with $M=500$ trajectories and $dt=0.02$.
All in all, we observe that noise generally leads to delocalization even in the presence of disorder for any collision rate $r_{c}\neq 0$. This behavior occurs because the noise-induced dephasing leads to progressive heating of the system, ultimately driving it toward a high-entropy $T=\infty$ steady state with the excitation uniformly distributed across all sites. Disorder slows down the process by creating localized regions that act as traps and inhibits transport by delaying the redistribution of energy. Consequently, while the system eventually reaches the $T=\infty$ state regardless of the disorder strength, the time required to do so increases with higher levels of disorder. This highlights the competing roles of noise, which promotes delocalization, and disorder, which instead favors localization.
In the case of collisions that are homogeneous in time and in space, that is e.g. with $r_{c}=0.1$ and $\nu=100$ in the figure, noise appears in well defined time slots. This makes visible the corresponding moments where magnetization and IPR change, exhibiting formation of plateaus in the IPR. In this case, we highlight that the higher the value of the disorder, the sharper the plateau. Also, we stress that the plateaus were not affected by the anisotropy. For $r_{c}<1$, the excitation exhibits localization, which persists until noise is introduced into the system at well-defined times (see Fig. 1(b)). These noise events induce delocalization, after which the excitation becomes localized again until a subsequent collision perturbs the system dynamics. This process repeats iteratively, leading the IPR to eventually approach its asymptotic value $1/N$, i.e. to the appearence of a fully delocalized state.
@@
In contrast, for $r_{c}\geq 1$, the frequency of collision events becomes sufficiently high to prevent any significant re-localization of the system. This behavior is particularly evident for $r_{c}=1$, as shown in Fig. 1(c). For $r_{c}\geq 1$, the system transitions to a regime dominated by frequent collisions, which suppress the transient localization observed at lower collision rates. In this regime, the IPR evolves smoothly over time without forming plateaus, indicating sustained delocalization. High collision rates prevent the system from stabilizing in a localized state, continuously redistributing the excitation across the chain. The comparison between homogeneous ($\nu\gg 1$) and heterogeneous ($\nu\leq 1$) noise regimes reveals distinct transport behaviors. In fact, for homogeneous noise, collisions occur at regular intervals, leading to predictable dephasing effects and smoother transitions between localization and delocalization. In contrast, heterogeneous noise introduces temporal variability in collisions, resulting in intermittent localization and a more irregular evolution of the IPR. This highlights the role of noise structure in modulating transport dynamics.

### v5

@@ Abstract
We systematically investigate the interplay of disorder and time-homogeneous collisional noise in shaping the transport dynamics of an anisotropic XXZ spin chain. Using stochastic collision models to simulate interaction with the environment, we explore the localization-delocalization transitions across regimes with single and multiple excitations. We find that space-homogeneous and low-rate collisions can shape regions where localization sets in the form of subsequent plateaus. The localization process has universal features for the plateaus duration and the delocalization time. Interactions among the excitations favor this process even for tiniest disorder levels.
@@ I Introduction
Quantum transport governs charge, spin and energy transfer in a wide variety of systems, from solid-state physics and spintronics to quantum technologies [YagoMalo2024, Ferry_2022, Brahlek2024, Sebolt2024]. In fact, by exploring the effects of disorder, noise, and interactions, we can gain insights into the mechanisms driving localization-delocalization transitions, thermalization, and the emergence of collective phenomena in low-dimensional and complex systems. In particular, current technologies enable the study of quantum transport with the engineering of reduced-dimensionality systems in solid state and in AMO (atomic, molecular, and optical) platforms [Wili2023, Huang2024].
@@
We are interested in transport of energy and excitations in biological systems [Plenio2008, SCM_approach_to_transport] and in coarse-grained models within the so-called quantum-like paradigm [YagoMalo2023, Khrennikov]. Regardless of the presence of quantum effects in biological processes, the open quantum systems framework [non_markov_review, Wiseman_Milburn, Petruccione_Breuer, Gardiner_Zoller] offers a well-suited mathematical toolbox for modeling complex phenomena, both quantum and classical [AlessandroetAl, Khrennikov]. Interestingly, open quantum networks modeled by paradigmatic many-body Hamiltonians can also be engineered in tunable current quantum simulators to reproduce essential transport dynamics [roy2024simulating].
@@
Quantum effects in biological systems present a fascinating yet debated area of research, challenging conventional views on the role of quantum mechanics in the warm and noisy environment of living organisms. These physics-based models can incorporate quantum effects into the description of microscopic phenomena, allowing us to study how a noisy and out-of-equilibrium biological environment is influenced by coherent and dissipative quantum effects [Plenio2008, OlayaCastro2008, pathania2024dynamics].
@@
In general, the behavior of transport in a quantum system serves as a powerful diagnostic tool for understanding thermalization in disordered conditions [dohner2023thermalization]. In the absence of interactions, disorder can induce Anderson localization [Guan2019], where transport is suppressed due to interference effects, preventing the system from reaching thermal equilibrium. Interactions or noise disrupts this localization mechanism, leading to a transition toward diffusive or even ballistic transport, depending on the system’s parameters [MBL_Nandkishore, MBL_from_IPR_study]. The transition from localized to delocalized regimes provides insight into how thermalization emerges in complex systems. Consequently, disorder is a mechanism that suppresses transport in isolated systems and, in open systems, also a factor that competes with noise to dictate the system’s relaxation dynamics.
@@
In [AlessandroetAl], we explored how dissipation modulates transport in an open quantum system using stochastic collision models accounting for space and time heterogeneous and homogeneous collisions. In so doing, we could focus on the interplay between noise, interactions, and coherence under sufficiently general conditions. By tuning noise parameters, we demonstrated controllable modifications to transport properties, identifying regimes where increasing the collision rate enhances transport. Thus, we showed that engineered dissipation can counteract localization effects.
@@
We now focus on the combination of collisional noise and disorder which provides a model that closely mirrors the dynamics of transport in many complex media including those relevant to biological systems. Collisional models replicate approximate interactions with noisy warm environment, while disorder introduces local spatial inhomogeneities. Together, these elements capture key features of energy transfer in systems where coherent and dissipative processes coexist, relevant in several fields of physics. Examples are charging and discharging performances in open quantum system batteries [Carrega2020, Mayo2022], transport in an engineered environment of complex systems [Nokkala2024] along with many body localization in 1-D systems [Levi2016]. Besides, the tunability of collisional noise parameters enables the simulation of biologically relevant scenarios, such as the exploitation of noise-assisted transport [Huelga2009, Huelga2013].
@@
In this paper, we investigate the transport properties and thermalization dynamics of a noisy, disordered XXZ spin chain [Prosen_Buca_correlations_XXZ, Prosen_Buca_XXZ, Rakov_Weyrauch_Braiorr-Orrs, McRoberts2024, KT_XXZ], modeled through stochastic collision processes. Among different possibilities, we adopt this many-body model for its capability of describing in a general manner tunneling-based transport of excitations in the presence of coherent anisotropic interactions. We identify localization/delocalization transitions and thus transport windows for multiple excitations tuning the interplay of disorder, anisotropy, and noise.
@@
The paper is organized as follows. In Section II we present the model of our disordered spin-chain interacting with the environment via stochastic collision models and introduce the figures of merit adopted to study transport properties. In Section III, we present the results of our study comparing the cases of one, two and multiple injected excitations also highlighting the entanglement properties in those regimes. Finally, in Section IV we present our conclusions and discuss future perspectives in the light of technological progress in quantum science and biology.
@@ II Model and methods
The conceptual model setting is sketched in Fig. 1 (a). We investigate an integrable generalization of the Heisenberg spin chain that accounts for uni-axial anisotropy [Franchini_chains], with the addition of a disorder term:
@@
$+\sum_{i=1}^{N}h_{i}\sigma^{z}_{i},$ (1)
and subjected to dissipation through noise in the form of stochastic collision models, as depicted in Sec. II.1, with open boundary conditions (OBC). In Eq. (1), $\sigma^{x/y/z}_{i}$ are the Pauli operators acting on spin located at site $i$; $J$ is the spin exchange rate, that can also be seen as the tunneling amplitude for excitation quasiparticles throughout the system; $\Delta$ governs the anisotropy in the form of a spin-spin interaction along the z-axis. The last term introduces disorder for every spin independently. This is in the form of a random local field, where each $h_{i}$ is drawn individually from the uniform distribution over the interval [h,h][-h,h]. In our analysis, we consider the regime with $J=1(>0)$ and then scale every other energy to $J$. As a result, the ground state with disorder $h=0$ is an anti-ferromagnet for $\Delta>1$, and a paramagnet for $0<\Delta<1$.
Figure 1: Sketch of our model. (a) Open quantum spin chain with energy levels affected by disorder in the [h,h][-h,h] range. Excitations are depicted as flipped spins (green). Their dynamics is affected by the tunneling rate $J$ (black arrows), interaction strength $\Delta$ (blue ovals) and the addition of a noisy environment. The latter is represented by the auxiliary light-brown qubits colliding with the sites over time. (b)-(c) Histograms displaying the number of collisions undergone over time tJ by the auxiliary qubits (log scale) against spins. This is shown for a noise shape parameter $\nu=100$ in the time-uniform regime, and collision rates $r_{c}=1$ (b) and (c) $r_{c}=0.1$. We use the same simulation parameters: number of sites in the spin is $N=41$ sites, $M=500$ trajectories and final time of simulation $tJ=30$.
When considering the time evolution of excitations or quench dynamics, the disorder term will control different transport regimes, ranging from ballistic to diffusive, and it can lead to the emergence of Many-Body localization (MBL) [Prosen2011]. In fact, disorder can lead to the breakdown of thermalization, meaning that the system fails to reach thermal equilibrium at long times resulting in transport suppression. Tuning the disorder through $h$ and analyzing the resulting transport behavior, it is possible to map the parameter regimes where localization or delocalization dominates.
@@ II.1 Noise modeling
So far we have discussed how the coherent and disorder-driven terms of our model are described. Here, we introduce dissipation through a particular discretization of the environment degrees of freedom (d.o.f.) that lies in the family of stochastic collision models [SCM_approach_to_transport, Gallina]. This type of environment description captures spatio-temporal homogeneity and heterogeneity in the system-bath interaction, effectively approximating biological environments [SCM_approach_to_transport, Cao2020]. As depicted in Fig. 1, the environment d.o.f are modeled as auxiliary qubits (light brown circles) that collide over time with the spins in the chain.
In this approach, we can govern the noise regimes via appropriate distributions describing the probability of system-environment interaction events. This results in a flexible environment description, including non-trivial temporal correlations even remaining within a Markovian-like description. As it is the case for many dissipative dynamical maps, stochastic collisional models are compatible with many local quantum channels. In our case, we account for density-density collisions decorrelating spins from the coherent evolution, thus effectively producing dephasing in the system degree of freedom.
We choose to describe the collisional rate between spins and auxiliary d.o.f. by adopting the especially flexible Weibull distribution [Weibull_Renewal_Processes, SCM_approach_to_transport]:
@@
with $\tau_{th}$ being the mean collision time and $\Gamma$ given by $\int_{0}^{\infty}\tau p(\tau)d\tau=\mu\Gamma(1+1/\nu)=t_{th}$ from Eq. (2)). In Fig. 1. (b)-(c), we present histograms for the number of collisions as a function of time for homogeneous noise ($\nu=100$) and two different collision rates modifying how recurrent is the interaction with the environment. Regarding the description of noise, we specify here that we use the same nomenclature as [SCM_approach_to_transport] where the temporal heterogeneity is defined by looking at the variance of the stochastic process and the regularity with which the collisions occur.
The collision times are generated independently for each site as follows: given a number taken randomly from the uniform distribution $u\in[0,1]$, by inverse-transforming sampling, the waiting time since the previous collision is going to be $t=\sqrt[\nu]{-\mu^{\nu}\ln\!\left|1-u\right|}$, with $\nu$ and $\mu$ being the shape and scale parameter of Eq. 2. The initial list of collision times $\bar{S}$ contains these first waiting times (interpreted as absolute times of the first collisions) for each site. After a collision occurs at site $i$ at time $S_{i}$, a new independent waiting time is drawn in the same way and added to $S_{i}$, so that $S_{i}\to S_{i}+t_{\text{new}}$. Thus, the next site in which a collision will occur will be simply given by the minimum of the updated list of waiting times $\bar{S}^{\prime}$.
When collisions occur according to the probability distribution, a quantum channel is applied [SCM_approach_to_transport, Decoherence_without_entanglement_Quantum_Darwinism]. This is given by:
@@
with $U_{\textrm{coll}}=\exp{-i(\pi/2)\sigma_{a}^{x}\otimes\sigma^{z}_{i}}$ representing the collision event and $t_{i}$ being the time before the given collision. Initially, the system is prepared in a well-defined state, which typically consists of one or more spins flipped along the chain. Subsequently, for each site of the system, a list of future collision times is generated according to a stochastic process governed by the chosen noise distribution. Between two consecutive collisions, the system evolves coherently, according to the dynamics dictated by the Hamiltonian in Eq.(1). When a collision occurs at a given site, the system undergoes a transformation described by the quantum channel in Eq.(4).
After the application of the quantum channel, the time for the next collision at the affected site is resampled from the same probability distribution. The entire procedure is then iteratively repeated, allowing for the simulation of the time evolution of the system, using exact diagonalization, in the presence of structured noise. For a detailed description, both of the noise model and the effects of the collisions, we refer to our previous work [AlessandroetAl]. We note here that we kept the same disorder realization across different values of the parameters, to isolate the consequences of changes due to collisions. We have verified that our conclusions remain unchanged under different disorder realizations.
@@ II.2 Characterizing transport
[…]
@@
Therefore, to characterize the emergence of parameter regions where the system (de)localizes, we also consider the behavior of the entanglement entropy for selected scenarios. The entanglement entropy constitutes one of the well-known heralds of MBL [MBL_Nandkishore, Jorge_Localization], as it presents logarithmic growth in many body localized systems. While not generally accessible in experiments, with a few important exceptions [Inn_toolbox, Greiner], the study of entanglement entropy in thermal and nonthermal many-body states [Banuls] has helped to understand and quantify the degree of correlation and information spreading between system parts. Its behavior can for example become non-trivial in regions close to the transition point [Banuls2]. The entanglement entropy can be characterized via the Von Neumann entropy [Nielsen_Chuang] defined as follows: consider a bi-partite system composed by $A$ and $B$ subsystems, then the entanglement entropy is $S_{vN}=\rho_{A}\log\rho_{A}$, with $\rho_{A}=Tr_{B}\rho$ the reduced density matrix of subsystem $A$. We are interested in regions in which transport can be modulated to occur only in well-defined time windows. Thus, we analyze the behavior of entanglement entropy as a complementary figure of merit.
@@ III.1 One excitation case
We first consider the single-excitation case, so to isolate the competition between quenched disorder (favoring localization [MBL_Nandkishore]) and noise (favoring delocalization), while avoiding complications from interactions between excitations.
We summarize in Fig. 2 the IPR behavior. Fig. 2(a) summarizes the long-time behavior of the Inverse Participation Ratio (IPR) at time $tJ=30$ as a function of the shape parameter $\nu$ and the collision rate $r_{c}$. We see that for collisions heterogeneous over time (low values of $\nu$) with a high rate of interaction (high $r_{c}$), the system exhibits higher localization (larger IPR) preventing the IPR to reach its asymptotic value in our observation window. This resembles the case of no disorder [AlessandroetAl], even if the localization is stronger here even for intermediate values of $r_{c}$. Instead when $\nu>>1$, the IPR approaches $1/N$. We now focus on Fig. 2(b)-(c) where we show the IPR evolution at fixed shape parameter $\nu=100$ in the regime of time-homogeneous collisions. The IPR time evolution is displayed in Fig. 2(b) for various collision rates $r_{c}$ and fixed large disorder range $h=10$. We then show in Fig. 2(c) the IPR versus time for different disorder range constants $h$ from 0.1 to 10 at fixed low collision rate $r_{c}=0.1$. In the case of collisions that are homogeneous in time and in space, that is e.g. with $r_{c}=0.1$ and $\nu=100$ in the figure, noise appears in well-defined time slots. This makes visible the corresponding moments where magnetization and IPR change, exhibiting formation of plateaus in the IPR. In this case, we highlight that the higher the value of the disorder, the sharper the plateau. The plateaus are observed for high values of the shape parameter, e.g. while collisions occur at regular intervals, since noise is introduced into the system at well-defined times (see Fig. 1(b)).
For $r_{c}<1$, localization followed by delocalization processes repeat iteratively for the excitation, leading the IPR to eventually approach its asymptotic value $1/N$, i.e. to the appearance of a fully delocalized state. We note that the cases with $\Delta\neq 0$ are not reported here because the inclusion of an anisotropic interaction term is seen to not qualitatively alter the system’s behavior in the single-excitation regime, as expected considering that it contributes a boundary effect. For more details see [AlessandroetAl].
For $r_{c}\geq 1$ instead, the system is driven towards a diffusive regime. In fact, in this case the system transitions to a regime dominated by frequent collisions, which suppress the transient localization observed at lower collision rates. In this regime, the IPR evolves smoothly over time without forming plateaus, indicating sustained delocalization. High collision rates prevent the system from stabilizing in a localized state, continuously redistributing the excitation across the chain. However, we can say that for $r_{c}>5$ the IPR increases with $r_{c}$, failing to reach his asymptotic value in our observation window. In fact, too high values of collision rate (in the figure $r_{c}=50,100$) slow down the dynamics and the spread of the excitation. This can be seen as a manifestation of the Zeno effect, as appeared also in the case of no disorder [AlessandroetAl]. This effect also causes the non-monotonic dependence on the collision rate which arises because collisions assist transport only when their rate is not too large. For $r_{c}\sim 5$ and above, frequent system-environment interactions act as repeated local projections that suppress coherent dynamics, slowing it, and finally freezing it, acting as strong disorder. The overlap between the curves in Fig. 2(b) for $r_{c}=0$ and $r_{c}=0.1$ is to be expected (see also Fig. 3(a) and Fig. 7): indeed, a high shape parameter $\nu=100$ causes the absence of collisions till $tJ=10$, thus the system can be considered effectively closed, up to that time.
@@
All in all, the comparison between homogeneous ($\nu\gg 1$) and heterogeneous ($\nu\leq 1$) noise regimes reveals distinct transport behaviors. In fact, for homogeneous noise, collisions occur at regular intervals, leading to predictable dephasing effects and smoother transitions between localization and delocalization. In contrast, heterogeneous noise introduces temporal variability in collisions, resulting in intermittent localization and a more irregular evolution of the IPR. This highlights the role of noise structure in modulating transport dynamics.
Figure 2: (De)localization behavior: the case of one excitation. Inverse Participation Ratio (IPR) for a spin chain with $N=41$ sites. (a) IPR at long time ($tJ=30$), for large disorder range ($h=10$) and different shape parameters $\nu$ and collision rates $r_{c}$. (b) IPR vs time for large disorder range ($h=10$) and different collision rates $r_{c}$. Curves with progressively increasing tickness and color scale refer to increasing values of $r_{c}=0,0.1,0.5,1,5,50,100$. (c) IPR vs time $tJ$ for low $r_{c}=0.1$ and different disorder range $h$. Notice the emergence of plateaus of temporary stabilization of localized behavior. In (b) and (c) the shape parameter is $\nu=100$ in the time-homogeneous collision regime. The data refer to the case of no anisotropy ($\Delta=0$ in Eq. (1)) since they are seen to be not significantly altered by finite $\Delta\neq 0$. The simulations were performed with $M=500$ trajectories and timestep timestep $dt=0.02$.
We observe that noise generally leads to delocalization even in the presence of disorder for any collision rate $r_{c}\neq 0$. This behavior occurs because the noise-induced dephasing leads to progressive heating of the system, ultimately driving it toward a high-entropy $T=\infty$ steady state with the excitation uniformly distributed across all sites. Disorder slows down the process by creating localized regions that act as traps and inhibits transport by delaying the redistribution of energy. Consequently, while the system eventually reaches the $T=\infty$ state regardless of the disorder strength, the time required to do so increases with higher levels of disorder. This highlights the competing roles of noise, which promotes delocalization, and disorder, which instead favors localization.